Compute all answers in (Solver) Excel, and show equations. Q2. Assume two markets; Pizza= Qs = 5+8Pp Qd = 10 - .2Pp + Ph Hamburger= Qs 20 + Ph Qd = 40 -3Ph + 2Pp Find the price of pizza and hamburger that clears both markets at the same time.
Answer #1

Answer : At market clearing price level, demand = supply occur.

For Pizza :

Qd = Qs

=> 10 - .2Pp + Ph =  5 + 8Pp

=> 10 - 5 + Ph = 8Pp + 0.2Pp

=> 5 + Ph = 8.2Pp

=> Pp = (5 + Ph) / 8.2

=> Pp = 0.61 + 0.12Ph

Now, for Hamburger :

Qd = Qs

=> 40 -3Ph + 2Pp = 20 + Ph

=> 40 - 3Ph + 2 (0.61 + 0.12Ph) = 20 + Ph [By putting the value of Pp]

=> 40 - 3Ph + 1.22 + 0.24Ph = 20 + Ph

=> 40 + 1.22 - 20 = Ph + 3Ph - 0.24Ph

=> 21.22 = 3.76Ph

=> Ph = 21.22 / 3.76

=> Ph = 5.64

By putting the value of Ph in Pp equation we get,

Pp = 0.61 + 0.12 * 5.64

=> Pp = 0.61 + 0.68

=> Pp = 1.29

Therefore, $1.29 clears the Pizza market and $5.64 clears the Hamburger market at the same time.